What is the 50th term of the fibonacci sequence?


1
Expert's answer
2021-09-27T12:27:39-0400

By the Binet's formula:

"f_{50}=\frac{1}{\sqrt{5}}[(1+\sqrt{5})^{50}-(1-\sqrt{5})^{50}]=12,586,269,025."